# Probing the limits of effective temperature consistency in actively driven systems

### v1

@@ III.2 Kinetic and potential energy partition
To probe deviations from equilibrium behavior, we naively define the tracer’s kinetic temperature as,
@@
This definition seems to result in a contradiction between the kinetic and potential temperatures, since deviations from the potential temperature $T_{\text{pot}}$ (Eq. 4) are observed for all $N_{b}$ configurations (solid gray line in Fig. 5A). Namely, $T_{\text{kin}}$ saturates and becomes largely independent of $N_{b}$ in the high-density limit. This clear mismatch is further supported by direct measurements of the effective trapping frequencies $\Omega$, extracted from the tracer’s velocity ACFs (see Fig. 3F). Across all system setups, $\Omega$ yields values lower than the natural frequency $\sqrt{k/m}$, even in an underdamped dynamic regime.
Figure 5: The modified kinetic temperature:  Data are presented for configurations with $N_{b}=\{3,4,6,10,12,15,17\}$ bbots, averaged over 375 one-minute tracer trajectories recorded in unperturbed conditions.  A. Comparison of the standard kinetic temperature $T_{\text{kin}}$ (solid gray line), calculated using the tracer mass $m\approx 1$ g (Eq. 5), the modified kinetic temperature $\tilde{T}_{\text{kin}}$ (solid red line), evaluated using Eq. 6, and the potential temperature $T_{\text{pot}}$ (dashed blue line). Error bars indicate standard deviations. Dot-dashed red lines show deviations in $\tilde{T}_{\text{kin}}$ arising from increasing the measurement time interval $\Delta t$.  B. Separate contributions to $\tilde{T}_{\text{kin}}$: (upper panel) the velocity variance $\sigma_{v}^{2}=\langle v_{x}^{2}\rangle$, and (lower panel) the effective mass $\tilde{m}$, evaluated from stationary ACF dynamics using Eqs. 1 and 6. Gray dot-dashed lines indicate deviations due to increasing $\Delta t$ in both panels.
However, the two effective temperatures agree, once the change in the effective mass is taken into consideration (solid red line in Fig. 5A). We define an effective mass, $\tilde{m}=\kappa\Omega^{-2}$, based on the velocity correlations. Similar mass corrections have appeared in prior studies of particles embedded in active or driven granular media, where persistent athermal fluctuations give rise to an apparent added mass effect [73, 74, 75]. In these cases, mapping the dynamics onto a linear Langevin equation, consistent with the ACF solution in Eq. 1, relies on short persistent times and typically requires a modification of system parameters [76, 77]. Thus, the kinetic temperature definition is modified to,
@@
where $\tilde{m}$ is defined as a $N_{b}$-specific effective mass, under a constant potential stiffness $\kappa$. Deviations of $\tilde{T}_{\text{kin}}$ measurements resulting from low measurement rates are shown as red dashed lines in Fig. 5A.
@@
Figure 6: Steady-state work FR test:  Results present averages over 375 one-minute perturbed tracer trajectories, under a strong external airflow applied by a fan at an operating voltage $13.5$ V. A. Work FR test for $N_{b}=10$. The temperature measurement of $T_{\text{FR}}$, extracted from a fit to Eq. 7, is consistent with the potential temperature ($T_{\text{pot}}$) determined from the unperturbed definition in Eq. 4.  B. Normalized plot of work FR tests across various $N_{b}=\{2,3,4,6,10,12,15,17\}$ configurations. The solid line indicates perfect agreement between the effective temperatures obtained from the work FR and from the equipartition definition, demonstrating thermodynamic-like consistency.
@@ III.3 Effective temperature based on work FR
Here we investigate a work FR in a NESS subjected to a constant force $F_{0}=\kappa\Delta x_{\epsilon}$ (applied at time $t\rightarrow$-$\infty$). An asymmetric work distribution is obtained by the energetic fluctuation around the steady-state, defined as $\delta W(t)=F_{0}\left[x_{\epsilon}(t)-\langle x_{\epsilon}\rangle\right]$. A steady-state work FR can then be expressed as [78, 79, 13],
@@
The dashed linear line represents $T_{\text{FR}}=T_{\text{pot}}$, where $T_{\text{pot}}$ is obtained independently in the unperturbed state (a). Namely, the two temperature measures agree, providing an experimental validation of the work FR (Eq. 7). This result can be extended to different $N_{b}$, as shown in Fig. 6B. Here, the solid linear line represents $T_{\text{FR}}=T_{\text{pot}}$. Deviations for large values of the work fluctuations are expected due to under-sampling. Clear violations of Eq. 7 are mainly observed for $N_{b}=3$ and $2$. Thus, for larger $N_{b}$ systems, $T_{\text{FR}}$ is also consistent with the FDR-based measure $T_{\text{eff}}$. Notably, these temperatures were measured in independent measurements under different perturbations.
@@ IV Limitations of a consistent effective temperature
To further assess the range of consistency of effective temperature measures, we repeated the experiments from Section 3 for other system configurations. The standard experimental setup (detailed in Fig. 1 and shown in Fig. 7A) was modified by varying both the passive tracer and the surrounding active particles. Specifically, these modifications included: using a smaller styrofoam tracer particle (Fig. 7B); employing faster active bbots (Fig. 7C); and introducing a heavy tracer particle (Fig. 7D), i.e., heavier than the bbots. Further system details are provided in the SI.
Figs. 7E,F compare the tracer’s modified kinetic temperature ($\tilde{T}_{\text{kin}}$) and potential temperature ($T_{\text{pot}}$) as a function of the number of bbots ($N_{b}$) for these configurations. The solid linear line represents the condition: $\tilde{T}_{\text{kin}}=T_{\text{pot}}$.
In the alternative experimental setups that include a light tracer (Figs. 7B,C), $\tilde{T}_{\text{kin}}$ and $T_{\text{pot}}$ coincide within measurement error and increase with $N_{b}$, consistent with our previous findings (Fig. 5). These systems can also yield consistent $T_{\text{eff}}$ measurements whether via FDR (Eq. 2) or work FR (Eq. 7) tests (see details SI). However, when replacing the lightweight tracers with a heavy tracer, we observe differences between the measured $T_{\text{pot}}$ and $\tilde{T}_{\text{kin}}$ (see Fig. 7F).
Ultra-high-speed videos (SM movie 2) reveal the mechanical origin of this deviation. For lightweight tracers, interactions with the self-propelled bbots occur as discrete impacts, yielding a measurable ballistic regime between collisions. In contrast, heavy tracers can experience rapid, persistent collisions with the bbots that occur on timescales shorter than the experimental sampling interval ($\Delta t$). Upon strong collisions the tracer visibly changes the bbot propulsion dynamics. In addition, the heavy tracer remains embedded among the bbots, whereas a lighter tracer may be kicked upwards, escaping entrapment.
@@ V Summary and discussion
In this work, we have experimentally demonstrated that a passive granular tracer, confined in a harmonic potential and driven by an active fluctuating medium, can sustain a NESS that is well-characterized by a single effective temperature $T_{\text{eff}}$. This effective temperature serves as a unifying parameter that encapsulates both the system’s dynamic response properties and its stationary fluctuations (as detailed in Fig. 2).
We identify three key features that appear essential to this equilibriumlike behavior of effective temperature consistency. First, both the spontaneous fluctuations of the tracer and its response to weak external perturbations are governed by the same physical mechanism, i.e. random collisions with the active medium [67]. This is manifested in the validity of a linear FDR [10, 60].  Second, the light tracer is only weakly coupled to the environment, which minimizes the chance of physical entrapment or caging by bath particles, an effect known to induce long-term memory in the tracer’s dynamics [80, 81].  Finally, the bath dynamics (bbot motion) are largely unaffected by the external driving (airflow) that acts mainly on the tracer particle, mimicking the timescale separation typical to equilibrium Brownian systems [39, 63, 66].
The agreement between different definitions of effective temperature observed here is unexpected for two reasons. First, the noise is a result of non-exponential collision times and the system exhibits non-trivial steady-state distributions, unlike the Poisson shot noise conditions leading to Boltzmann-like steady-states under which such agreement has been predicted [82, 26, 83]. Second, the use of an inertial tracer introduces memory effects in active matter, often breaking Markovian dynamics and precluding a single consistent effective temperature [84, 81, 85, 66]. That this agreement persists despite these deviations from the idealized theoretical framework suggests a broader robustness of effective temperature concepts than previously recognized.
@@ Acknowledgements
Additional results and measurements
We furthermore provide the following supplemental material:  FDR tests under constant $N_{b}$ and increasing perturbation amplitudes, i.e., fan operating voltage (Fig. 10);  Individual FDR results for position and velocity tracer observables (Figs. 11,12), and work FR results (Fig. 13), with different $N_{b}$. The tracer’s restitution coefficient measurement (Fig. 13); Particularly, these results were obtained with the standard experimental setup, as detailed in the main paper.
Fig. 15 and 16 present FDR and FR tests with alternative system configurations. Namely, we employ faster bbots (15) and a smaller styrofoam ball as a passive tracer (16). The position autocorrelation functions are approximately proportional to the responses with $T_{\text{eff}}=T_{\text{pot}}$, yet the FDR in the velocity observable show violations for short times. The work FR is valid for sufficiently small values, from which a consistent effective temperature can be extracted with $T_{\text{FR}}\approx T_{\text{eff}}$. We note that reliable $T_{\text{eff}}$ measurements across different $N_{b}$ would require additional tuning of the perturbation amplitude and the potential stiffness, to ensure stable dynamics in the system. The FDR tests generally fail for all considered setups that employ a heavy tracer under the air-driven perturbation (Fig. 17).
@@
[…]
Figure 14: Tracer’s coefficient of restitution versus impact velocities. The restitution coefficient $r$ was determined from the sound signal emitted by the styrofoam ball (the tracer) bouncing repeatedly off a surface, as detailed in Ref. [86]. We obtain $r$ by measuring the time lag between consecutive impacts: $r_{i}=|v^{\prime}_{i}/v_{i}|=(t_{i+1}-t_{i})/(t_{i}-t_{i-1})$, for an impact velocity $v_{i}=\frac{g}{2}(t_{i-1}-t_{i})$, $i=2,3,$.... Here, $v_{i}$ and $v^{\prime}_{i}$ are the normal pre- and post-collision velocities occurring at time $t_{i}$, and $g=9.81\textrm{m/s}^{2}$ is the gravitational acceleration. The plot shows four successive impacts of a repeated experiment (50 trials), with mean values from last to first respectively: $r=0.79$ (circle), $0.78$ (diamond), $0.76$ (square), $0.74$ (triangle), for $v_{\perp}=68.37$, $87.35$, $113.4$, $152.83$ cm/s.
Figure 15: Fast bbot setup.  A. The active bath contains $N_{b}=6$ fast bbots, the experimental setup is otherwise the same as in Figs. 10-13.  B. Average individual speed of a single bbot walking in a parabolic arena. Standard bbots (circle) and fast bbots (diamond), as referred to in the main paper.  C. FDR test for the tracer’s position and velocity observables, under the same perturbation setup as detailed in Figs. 11,12. The effective temperature is $T_{\text{eff}}=\kappa\langle x^{2}\rangle_{0}$.  D. Work FR test as detailed in Fig. 13.
Figure 16: Small tracer setup.  A. A small styrofoam ball is used as a passive tracer (33 cm diameter, $0.8$ g). The active bath contains $N_{b}=10$ standard bbots.  B. FDR test for the tracer’s position and velocity observables, under the same perturbation setup as detailed in Figs. 11,12. The effective temperature is $T_{\text{eff}}=\kappa\langle x^{2}\rangle_{0}$.  C. Work FR test as detailed in Fig. 13.
Figure 17: Heavy tracer setup.  A. A heavy plastic ball is used as a passive tracer (33 cm diameter, $8.2$ g). The active bath contains $N_{b}=10$ standard bbots.  B. FDR test for the tracer’s position and velocity observables, under the same perturbation setup as detailed in Figs. 11,12. A strong airflow ($13.5$ V fan operating voltage) was used as a step-perturbation. The effective temperature is $T_{\text{eff}}=\kappa\langle x^{2}\rangle_{0}$. The correlation functions fail to capture the response to the step-perturbation.

### v2

@@ III.2 Kinetic and potential energy partition
To probe deviations from equilibrium behavior, we define the tracer’s kinetic temperature as,
@@
As seen in Fig. 5, this definition seems to result in a contradiction between the kinetic and potential temperatures, since deviations from the potential temperature $T_{\text{pot}}$ (Eq. 4) are observed for all $N_{b}$ configurations (solid gray line in Fig. 5A). Namely, $T_{\text{kin}}$ saturates and becomes largely independent of $N_{b}$ in the high-density limit. This clear mismatch is further supported by direct measurements of the effective trapping frequencies $\Omega$, extracted from the tracer’s velocity ACFs (see Fig. 3F). Across all system setups, $\Omega$ yields values lower than the natural frequency $\sqrt{\kappa/m}\approx 5.3~\rm{s^{-1}}$, even in an underdamped dynamic regime.
Figure 6: Steady state work FR test: Results present averages over 375 one-minute perturbed tracer trajectories, under a strong external airflow applied by a fan at an operating voltage $13.5$ V. A. Work FR test for $N_{b}=10$. The upper panel shows an asymmetric work ($\delta W$) distribution in a NESS under a strong continuous perturbation (airflow). The temperature measurement of $T_{\text{FR}}$, extracted from a fit to Eq. 7 (small $\delta W$, solid line), is consistent with the potential temperature ($T_{\text{pot}}$, dashed line) determined from the unperturbed definition in Eq. 4. B. Normalized plot of work FR tests across various $N_{b}=\{2,3,4,6,10,12,15,17\}$ configurations. The solid (black) line indicates an ideal agreement between the effective temperatures obtained from the work FR and from the equipartition definition ($T_{\text{pot}}=T_{\text{FR}}$), demonstrating thermodynamic-like consistency.
We note, however, that for tracer particles embedded in active or driven granular media, an effective mass emerges as a consequence of persistent athermal fluctuations and memory effects [51, 71, 33]. Accordingly, we define an effective mass, $\tilde{m}=\kappa\Omega^{-2}$, derived from the velocity correlations. Mapping the dynamics onto a linear Langevin equation, consistent with the ACF solution in Eq. 1, relies on short persistence times and typically necessitates a modification of system parameters [85, 73]. Consequently, the definition of the kinetic temperature is modified to,
@@
where $\tilde{m}$ is defined as a $N_{b}$-specific effective mass, under a constant potential stiffness $\kappa$.
Using the effective mass to compute the kinetic temperature yields agreement between the two effective temperatures (solid red line in Fig. 5A). We note that the effective mass correction is derived from the correlation function of the unperturbed NESS, independent of both the step-response measurement and the FDR result (previous section). Deviations of $\tilde{T}_{\text{kin}}$ measurements resulting from low measurement rates are shown as red dashed lines in Fig. 5A.
@@ III.3 Effective temperature based on work FR
Here we investigate a work FR in a NESS subjected to a constant force $F_{0}=\kappa\Delta x_{\epsilon}$ (applied at time $t\rightarrow$-$\infty$). An asymmetric work distribution is obtained by the energetic fluctuation around the steady state, defined as $\delta W(t)=F_{0}\left[x_{\epsilon}(t)-\langle x_{\epsilon}\rangle\right]$. A steady state work FR can then be expressed as [82, 58, 36],
@@
The dashed linear line represents $T_{\text{FR}}=T_{\text{pot}}$, where $T_{\text{pot}}$ is obtained independently in the unperturbed state (a). Namely, the two temperature measures agree, providing an experimental validation of the work FR (Eq. 7). This result can be extended to different $N_{b}$, as shown in Fig. 6B. Here, the solid linear line represents $T_{\text{FR}}=T_{\text{pot}}$. Deviations for large values of the work fluctuations are expected due to under-sampling. Clear violations of Eq. 7 are mainly observed for $N_{b}=2$ - $4$. For larger $N_{b}$ systems, $T_{\text{FR}}$ is also consistent with the FDR-based measure $T_{\text{eff}}$. Notably, these temperatures were measured in independent measurements under different perturbations. This result shows that, for a broad range of parameters ($N_{b}=6$ - $17$), the same effective temperature that governs static (equipartition) and dynamic (FDR) properties of the system also rules the irreversibility of energy exchanges.
@@ IV Limitations of a consistent effective temperature
To further assess the range of consistency of effective temperature measures, we repeated the experiments from Section 3 for other system configurations. The standard experimental setup (detailed in Fig. 1 and shown in Fig. 7A) was modified by varying both the passive tracer and the surrounding active particles. Specifically, these modifications included: using a smaller styrofoam tracer particle (Fig. 7B); employing faster active bbots (Fig. 7C); and introducing a heavy tracer particle (Fig. 7D), i.e., heavier than the bbots ($m_{\text{bot}}=7.1$ g). Further system details are provided in the SI (see figures 6-8 therein).
Figs. 7E,F compare the tracer’s modified kinetic temperature ($\tilde{T}_{\text{kin}}$) and potential temperature ($T_{\text{pot}}$) as a function of the number of bbots ($N_{b}$) for these configurations. The solid linear line represents the condition: $\tilde{T}_{\text{kin}}=T_{\text{pot}}$. In the alternative experimental setups that include a light tracer (Figs. 7B,C), $\tilde{T}_{\text{kin}}$ and $T_{\text{pot}}$ coincide within measurement error and increase with $N_{b}$, consistent with our previous findings (Fig. 5). These systems can also yield consistent $T_{\text{eff}}$ measurements whether via FDR (Eq. 2) or work FR (Eq. 7) tests (see details in SI). However, when replacing the lightweight tracers with a heavy tracer, we observe differences between the measured $T_{\text{pot}}$ and $\tilde{T}_{\text{kin}}$ (see Fig. 7F).
Ultra-high-speed videos (SM movie 2) reveal the mechanical origin of this deviation. For lightweight tracers, interactions with the self-propelled bbots occur as discrete impacts, yielding a measurable ballistic regime between collisions. In contrast, heavy tracers can experience rapid recurrent collisions with the bbots, occurring on timescales shorter than the experimental sampling interval ($\Delta t$). Upon strong collisions the tracer visibly changes the bbot propulsion dynamics. In addition, the heavy tracer remains embedded among the bbots, whereas a lighter tracer may be kicked upwards, escaping entrapment.
@@ V Summary and discussion
In this work, we have experimentally demonstrated that a passive granular tracer, confined in a harmonic potential and driven by an active fluctuating medium, can sustain a NESS that is well-characterized by a single effective temperature $T_{\text{eff}}$. This effective temperature serves as a unifying parameter that encapsulates both the system’s dynamic response properties and its stationary fluctuations (as detailed in Fig. 2). We find that introducing an effective mass is essential to properly define the tracer’s kinetic temperature, ensuring its consistency with the effective temperature of the NESS.
We identify three key features that appear essential to this equilibriumlike behavior of effective temperature consistency. First, both the spontaneous fluctuations of the tracer and its response to weak external perturbations are governed by the same physical mechanism, i.e. random collisions with the active medium [7]. This is manifested in the validity of a linear FDR [47, 4]. Second, the light tracer is only weakly coupled to the environment, which minimizes the chance of physical entrapment or caging by bath particles, an effect known to induce long-term memory in the tracer’s dynamics [6, 13]. Finally, the bath dynamics (bbot motion) and the fluctuations they induce are largely unaffected by the external driving (airflow) that acts mainly on the tracer particle [34, 67, 40].
The agreement between different definitions of effective temperature observed here is unexpected for two reasons. First, the noise is a result of non-exponential collision times with the bath particles that show temporal correlation due to active self-propulsion. Under these conditions, the system exhibits non-trivial steady state distributions, in variance with the Poisson shot noise conditions leading to Boltzmann-like steady states under which such agreement has been predicted [22, 20, 45]. Second, the use of an inertial tracer introduces memory effects in active matter, often breaking Markovian dynamics and precluding a single consistent effective temperature [12, 13, 77, 40]. That this agreement persists despite these deviations from the idealized theoretical framework suggests a broader robustness of effective temperature concepts than previously recognized.
@@ Acknowledgements
Further results and measurements
We provide the following supplemental material: The average force exerted by the fan on the tracer was measured at different distances under a constant air stream (without bbots). The variance and skewness of the tracer’s stationary distributions (with $N_{b}=10$) are presented for different air stream amplitudes, corresponding to varying fan voltages (Fig. 1); FDR tests under constant $N_{b}$ and increasing perturbation amplitudes (Fig. 2); Individual FDR results for position and velocity tracer observables (Figs. 3,4) and work FR results (Fig. 5), with different $N_{b}$. Particularly, these results were obtained with the standard experimental setup, as detailed in the main paper.
Fig. 6 and 7 present FDR and FR tests with alternative system configurations. Namely, we employ faster bbots and a smaller styrofoam ball as a passive tracer, respectively. The position autocorrelation functions are approximately proportional to the responses with $T_{\text{eff}}=T_{\text{pot}}$, yet the FDR in the velocity observable show violations for short times. The work FR is valid for sufficiently small values, from which a consistent effective temperature can be extracted with $T_{\text{FR}}\approx T_{\text{eff}}$. These measurements fail with a heavy tracer setup (Fig. 8), as discussed in the main text.
We note that additional reliable $T_{\text{eff}}$ measurements across different $N_{b}$ would require additional tuning of the perturbation amplitude and the potential stiffness, to ensure stable dynamics in the system.
Figure 1: Forces and stationary statistics under external air stream. A, B. Average forces acting on the tracer in the absence of bbots, on a flat surface. Measurements were obtained using a force sensor (Mark-10 Force Gauge) with 1010V operating fan voltage: (A) as a function of xyx-y position; (B) along the main air stream axis $x$ (lower panel) - and as a function of height $z$ (upper panel), at a distance x=8x=8cm marked by the vertical dashed line. Error bars are standard deviations of $50$ measurements. C. Stationary probability distributions of the tracer’s position under different air stream intensities in the harmonic trap (unperturbed, 1010V, and 13.513.5V fan operating voltage), with $N_{b}=10$. D. Corresponding position variances. Results were obtained from time and ensemble averages over $M=375$ realizations of one-minute tracer trajectories. Error bars indicate standard deviations.
Figure 5: Work FR results in (strongly) perturbed steady-states with different numbers of bbots. A,B. Work FRs with $N_{b}=\{2,3,4,6,10,12,15,17\}$. The solid line represent a linear fit with a slope of $\sim 1/T_{\text{FR}}$. The dashed line has a slope of $\sim 1/T_{\text{pot}}$. The results were obtained for an ensemble average of $M=375$ perturbation sequences, with the same tracer $m\approx 1$ g, gravitational stiffness $k\approx 28.2~\text{g/s}^{-2}$, and fan operating voltage of $13.5$ V.
Figure 6: Fast bbot setup. A. The active bath contains $N_{b}=6$ fast bbots, the experimental setup is otherwise the same as in Figs. 2-7. B. Average individual speed of a single bbot walking in a parabolic arena. Standard bbots (circle) and fast bbots (diamond), as referred to in the main paper. C. FDR test for the tracer’s position and velocity observables, under the same perturbation setup as detailed in Figs. 3,4. The effective temperature is $T_{\text{eff}}=\kappa\langle x^{2}\rangle_{0}$. D. Work FR test as detailed in Fig. 5.
Figure 7: Small tracer setup. A. A small styrofoam ball is used as a passive tracer (33 cm diameter, $0.8$ g). The active bath contains $N_{b}=10$ standard bbots. B. FDR test for the tracer’s position and velocity observables, under the same perturbation setup as detailed in Figs. 3,4. The effective temperature is $T_{\text{eff}}=\kappa\langle x^{2}\rangle_{0}$. C. Work FR test as detailed in Fig. 5.
Figure 8: Heavy tracer setup. A. A heavy plastic ball is used as a passive tracer (33 cm diameter, $8.2$ g). The active bath contains $N_{b}=10$ standard bbots. B. FDR test for the tracer’s position and velocity observables, under the same perturbation setup as detailed in Figs. 3,4. A strong airflow ($13.5$ V fan operating voltage) was used as a step-perturbation. The effective temperature is $T_{\text{eff}}=\kappa\langle x^{2}\rangle_{0}$. The correlation functions fail to capture the response to the step-perturbation. C. Work FR test as detailed in Fig. 5, does not yield an effective temperature.